In an experiment carried at 1atm, a solution of 90.3 g of non-volatile compound soluble in
200 mL water boils at 101.451 o
C. What is the molar mass of the unknown compound?
Assume boiling point of water is 100.0 o
C, density of water 1.000 g cm-3
.
Kb = 0.512 o
C m-1
1
Expert's answer
2020-12-14T14:49:16-0500

"Tb=i\u00d7kb\u00d7m"

101.451 =2× 0.512 ×m

m =(101.451/1.024) =

m= 99.07

Molar mass = 99.07